# Long-term foehn reconstruction combining unsupervised and supervised learning Reto Stauffer, Achim Zeileis, Georg J. Mayr

### v1

@@ 2.2 ERA5 reanalysis
Figure 2 shows the target locations (C
; center) and
@@
geographically relative to the target station upstream (U
) and
downstream (D
) of the main foehn wind direction as well as to the right
(R
) and left (L
) of it. While the interpolated information from
the target location itself (C
) is always used as possible covariate for
@@ 3.1 Unsupervised learning: Mixture model for foehn classification
implemented in the R package foehnix
(Stauffer, 2023).
@@ 3.4 Season-trend decomposition
via the R package stR
(Dokumentov and Hyndman, 2023).
@@ Appendix A ERA5 fields
orthogonal from the center location (station location; C
) to the
@@
points upwind of the foehn wind direction (U
, UU
) and two
points in the downwind direction (D
, DD
) are defined with a
@@
side (*R
) and left hand side (*L
) of the main foehn direction
in the upwind direction (UR
, UL
) and downwind direction
(DR
, DL
). The downwind direction for stations located north
@@
C
toU
,U
toC
,UL
toDR
and others, -
@@
C
toU
,U
toC
,UL
toDR
and others, -
@@
C
toU
,U
toC
,UL
toDR
and others, -
@@
U
,UU
,UR
,UL
) as well as on the downwind side (sum ofD
,DD
,DR
,DL
) as well and the difference between these two sums, -
@@
U
,UU
,UR
,UL
) as well as on the downwind side (sum ofD
,DD
,DR
,DL
) as well and the difference between these two means.
@@ Appendix C Supervised learning: Reconstruction
the R package xgboost
(Chen et al., 2023) with
@@
eta
, learning rate: -
max_depth
, maximum depth of each weak learner (tree): -
min_child_weight
, minimum sum of instance weight needed in a child: -
gamma
, minimum loss reduction required:

### v2

@@ 2.2 ERA5 reanalysis
Figure 2 shows the target locations (C; center) and
@@
geographically relative to the target station upstream (U) and
downstream (D) of the main foehn wind direction as well as to the right
(R) and left (L) of it. While the interpolated information from
the target location itself (C) is always used as possible covariate for
@@ 3.1 Unsupervised learning: Mixture model for foehn classification
implemented in the R package foehnix (Stauffer, 2023).
@@ 3.4 Season-trend decomposition
via the R package stR (Dokumentov and Hyndman, 2023).
@@ Appendix A ERA5 fields
orthogonal from the center location (station location; C) to the
@@
points upwind of the foehn wind direction (U, UU) and two
points in the downwind direction (D, DD) are defined with a
@@
side (*R) and left hand side (*L) of the main foehn direction
in the upwind direction (UR, UL) and downwind direction
(DR, DL). The downwind direction for stations located north
@@
CtoU,UtoC,ULtoDRand others, -
@@
CtoU,UtoC,ULtoDRand others, -
@@
CtoU,UtoC,ULtoDRand others, -
@@
U,UU,UR,UL) as well as on the downwind side (sum ofD,DD,DR,DL) as well and the difference between these two sums, -
@@
U,UU,UR,UL) as well as on the downwind side (sum ofD,DD,DR,DL) as well and the difference between these two means.
@@ Appendix C Supervised learning: Reconstruction
the R package xgboost (Chen et al., 2023) with
@@
eta, learning rate: -
max_depth, maximum depth of each weak learner (tree): -
min_child_weight, minimum sum of instance weight needed in a child: -
gamma, minimum loss reduction required: